# A POD-TANN approach for the multiscale modeling of materials and macroelement derivation in geomechanics

### v1

@@ 5 Discussion and conclusions
Despite the linearity of the dimensionality reduction method, the obtained compression ratios, defined as $CR=1-\text{dim}(\boldsymbol{\mathcal{Z}})/\text{dim}(\boldsymbol{\xi})\%$, are very satisfactory, always higher than $90\%$. It is noted that for the case of the RUCs with ellipsoidal and leaf-shaped inclusions, IC sets of around 100000 degrees of freedom are considered, resulting in an ISV vector of only 25 components and $CR$ equal to $99.97\%$. The high compression ratio obtained allows for lightening the hidden layers of the TANN energy network. For the considered examples, a network with one hidden layer of 100 neurons is used. For most of the cases, the training is achieved fast after 1000 epochs of optimization. The whole procedure takes around five minutes of running time on common 64-bit machines with 4 cores. Therefore, the method is accurate but also very efficient, and it can be used as a handy solution for the homogenization of composite materials as well as a preliminary analysis tool when the materials of interest present a more complex microstructure. The stress predictions obtained using the proposed methodology have a mean absolute error of the order of 1e-4.
@@
Additionally, throughout the paper, a loss function that does not include energy terms was considered for training the energy network, see Section 3.3. The effectiveness of this function is demonstrated in the application section. The absence of energy terms extends the application of the procedure to data from different sources, such as experimental data, where energy measurements are difficult to obtain.
@@
A further benefit of the proposed procedure is its extensibility. The dimensionality reduction obtained through POD is linear and, in cases of strong nonlinearities in the behavior of the microstructure, might not permit obtaining a small-dimensional vector of ISV at the macroscale (irreducibility). Nonlinear dimensionality reduction methods can be used to further encode the POD reduced set (see [38] and [39] for the application of POD in series with nonlinear autoencoders and SINDy [40] or [41] for the use of multiple local reduced-order bases (ROBs) to tackle non-reducible problems by splitting them into multiple reducible ones with the aid of physics-informed cluster analysis).
@@
The developed technique shows great promise for the reconstruction task, particularly in reference to the reconstructed structural field in the macroelement problem. The ability to accurately reconstruct the microscopic fields from the macroscopic ISVs, as demonstrated in the monopile example, underscores the potential of the methodology for practical engineering applications. The combination of POD and TANN provides a robust and efficient framework for both homogenization and localization, enhancing the utility of the approach for complex geotechnical systems.

### v2

@@ 5 Discussion and conclusions
Throughout the paper, a loss function that excludes energy terms was used for training the energy network (see Section 3.3). The effectiveness of this function is demonstrated in the application section. By omitting energy terms, the procedure becomes applicable to data from various sources, such as experimental data, where energy measurements are challenging to obtain.
Despite the linear nature of the dimensionality reduction method, the resulting compression ratios - defined as defined as $CR=1-\text{dim}(\boldsymbol{\mathcal{Z}})/\text{dim}(\boldsymbol{\xi})\%$ are highly satisfactory, consistently exceeding 90%. Notably, for RUCs with ellipsoidal and leaf-shaped inclusions, IC sets encompassing approximately 100000 degrees of freedom are reduced to an ISV vector with only 25 components, yielding a $CR$ of 99.97%. This high compression ratio significantly reduces the complexity of the hidden layers in the TANN energy network. In the examples considered, a network with a single hidden layer of 100 neurons was sufficient. In most cases, training converged rapidly, typically after 1000 epochs of optimization. The entire procedure requires approximately five minutes of runtime on standard 64-bit machines with 4 cores. Thus, the method is not only accurate but also highly efficient. The macroscopic stress predictions generated by this methodology have a mean absolute error on the order of 1e-4.
The proposed POD-TANN approach may also serve as a valuable preliminary analysis tool for materials with more complex microstructures. This is particularly relevant for materials exhibiting strong nonlinearities, such as those undergoing softening, where localization phenomena within the RUC can take place. In such cases, classical homogenization techniques, like Asymptotic Expansion Homogenization, typically fail to account for bifurcation or localization phenomena, limiting their applicability in the presence of these complex behaviors.
The challenge with strong nonlinearities lies in the absence of a general homogenization framework that can accurately handle such cases. While AEH provides a rigorous approach for homogenizing periodic microstructures, its formulation does not allow the inclusion of bifurcation phenomena. This limitation prevents its direct application in scenarios where localization occurs, as it is inherently restricted to smooth solutions within a standard continuum framework.
However, the TANN formulation, as presented here, is not tied exclusively to AEH or any specific homogenization technique. This opens the door for alternative frameworks to be considered. Importantly, the TANN methodology for homogeneous materials is not restricted to Cauchy continua, as demonstrated in previous work, [7]. Therefore, if a robust homogenization framework capable of accounting for bifurcation phenomena were developed, particularly one based on higher-order continuum theories, the proposed POD-TANN workflow could be applied.
Such a framework would enable the TANN to model more complex microstructures and material behaviors, including those influenced by localization and bifurcation, thereby broadening the scope of the approach.
5.1 Frontiers on the reconstruction task
An important aspect in the reconstruction of microscopic fields is the fulfillment of boundary conditions.
In the presented examples on periodic RUCs, the POD-TANN approach, while effective in reducing the dimensionality of the system and training efficient models, does not inherently guarantee the enforcement of periodic BCs in decoding. This limitation can introduce reconstruction errors, which depend on the quantity of interest and the various sources of error within the process.
More specifically, in cases where the primary goal is to reconstruct the strain field, the error in fulfilling periodic BCs tends to be directly related to the number of retained POD modes. Since the macroscopic strain is an input to the TANN (for strain-driven formulations), any mismatch between the predicted macroscopic strain and the microscopic strain field reconstruction will affect the overall periodicity. As such, the error will generally decrease as more POD modes are incorporated, ensuring a more accurate representation of the microscopic behavior. This relationship suggests the possibility of developing a POD mode selection criterion explicitly based on the error associated with periodic BC reconstruction. This could complement or even extend the energy-based selection criterion currently used for internal state variable (ISV) selection, allowing for a more tailored approach depending on whether accuracy in boundary condition fulfillment is prioritized. This goes to direction of applying PINNs losses for the resolution of BVPs, see, e.g., [43] or [44].
A key consideration here is that the reconstruction of microscopic stress fields is not directly governed by the macroscopic stress predictions from the TANN, as no explicit localization operator maps the macroscopic stress to its microscopic counterpart. Instead, the microscopic stress field is indirectly reconstructed through the strain field and the internal coordinates. Once the microscopic strain field and ISVs are accurately reconstructed, the stress can be computed using the microscopic constitutive laws of the material (always known in numerical examples, but possibly unknown for more general scenarios). This sequential process introduces additional error sources, particularly when periodic boundary conditions are involved, as any inaccuracy in the strain field reconstruction could propagate into the stress field. Therefore, ensuring the periodicity of the reconstructed strain field becomes crucial for minimizing errors in the subsequent stress field calculations.
Moreover, the complexity of accurately reconstructing the periodicity in fields such as strain or stress depends not only on the number of POD modes but also on the nature of the microstructure and the loading conditions. In systems with strong heterogeneities or non-linear behaviors, this error may grow, necessitating more sophisticated selection strategies or even hybrid approaches that combine linear reduction methods (such as POD) with non-linear techniques to capture more complex interactions at the micro-scale. By linking the number of POD modes directly to the fulfillment of boundary conditions, this could allow for a more nuanced trade-off between computational efficiency and reconstruction accuracy, especially in scenarios where enforcing periodicity is critical to the validity of the simulation.
@@ 5.1 Frontiers on the reconstruction task
In this view, a further advantage of the proposed procedure is its extensibility. The dimensionality reduction achieved through POD is linear and, in cases where the microstructure exhibits strong nonlinearities, may not yield a sufficiently small-dimensional ISV vector at the macroscale (irreducibility). To address this, nonlinear dimensionality reduction methods can be employed to further encode the POD-reduced set. For instance, POD can be used in conjunction with nonlinear autoencoders and SINDy, as demonstrated in [45] and [46]. Additionally, multiple local reduced-order bases (ROBs) can be utilized to manage non-reducible problems by decomposing them into multiple reducible ones, aided by physics-informed cluster analysis (see [47] and [48]).
5.1.1 Frontiers on the macroelement derivation and use
@@
The developed technique demonstrates significant potential in reconstructing the structural field, particularly in the context of macroelement derivation. The ability to accurately reconstruct microscopic fields from macroscopic ISVs, as shown in the monopile example, underscores the practical utility of this methodology for engineering applications. By integrating POD and TANN, a robust and efficient framework is provided for both homogenization and localization, enhancing its applicability to complex geotechnical systems.
However, as highlighted in Section 4.3.4, the ”re-usability” of macroelements is currently limited due to the specificity of the computational setup, which is tailored to individual scenarios. This restriction necessitates retraining and reconfiguration when adapting to new conditions, such as changes in geometry, material properties, or boundary conditions. In this regard, Transfer Learning (TL) emerges as a promising avenue to overcome these limitations. By enabling the fine-tuning of pre-trained networks with smaller datasets, TL could extend the adaptability of macroelements across different scenarios, reducing the computational effort and enhancing the method’s scalability.
While the current framework is already effective for specific cases, the integration of TL presents a frontier for further expanding the flexibility and efficiency of the macroelement derivation process. This could enable the use of a generalized parametric macroelement model capable of adjusting to varying conditions, thereby broadening the application scope of the POD-TANN approach in geotechnical engineering.
@@ 6 Acknowledgment
Appendix A: Dimensionality reduction of ICs via POD
Mathematically, the POD allows to write a generic field $\boldsymbol{a}(\boldsymbol{x},t)$ by means of a modal expansion (see, e.g., Lumley [10]):
in which the column of $\boldsymbol{\Phi}$, $\boldsymbol{\Phi}_{k}\left(\boldsymbol{x}\right)$, are deterministic time-independent spatial functions (POD modes), modulated by the time-dependent coefficients $c_{k}(t)$, collected in the column vector $\boldsymbol{c}(t)$. The data-driven problem associated to the use of POD, is to find the optimal basis of modes, $\boldsymbol{\Phi}$, from data of $\boldsymbol{a}$. In practice, this is achieved exploiting the Singular Value Decomposition (SVD), [12].
In the following, $\boldsymbol{A}$ is used to denote a generic snapshot matrix, i.e., a collection of column vectors of snapshots of the fields $\boldsymbol{a}(\boldsymbol{x},t)$, at subsequent time instants. The notation in equation 19 is thus transformed into $\boldsymbol{A}=\boldsymbol{\Phi}\boldsymbol{C}$, where $\boldsymbol{C}$ collects the time varying coefficients of each mode, each column associated to a different time instant.
The SVD is a unique matrix decomposition that exists for every complex valued matrix $\boldsymbol{A}\in\mathbb{C}^{n\times m}$, [12]:
$\boldsymbol{A}=\boldsymbol{U}\boldsymbol{S}\boldsymbol{V}^{*},$ (20)
where $\boldsymbol{U}\in\mathbb{C}^{n\times n}$ and $\boldsymbol{V}\in\mathbb{C}^{m\times m}$ are unitary matrices with orthonormal columns, and $\boldsymbol{S}\in\mathbb{R}^{n\times m}$ is a matrix with real non-negative entries in the diagonal and zeros elsewhere, with ()(*)( ∗ ) the complex conjugate transpose.
When $n>m$, $\boldsymbol{S}$ has at most $m$ non-zero elements. The following exact economy representation of $\boldsymbol{A}$ is therefore possible:
$\boldsymbol{A}=\left[\boldsymbol{\hat{U}}\quad\boldsymbol{\hat{U}}^{\perp}\right]\left[\boldsymbol{\hat{S}}\quad\boldsymbol{0}\right]^{T}\boldsymbol{V}^{*}=\boldsymbol{\hat{U}}\boldsymbol{\hat{S}}\boldsymbol{V}^{*}.$ (21)
The columns of $\boldsymbol{\hat{U}}^{\perp}$ span a vector space that is complementary and orthogonal to that spanned by $\boldsymbol{\hat{U}}$. The diagonal elements of $\boldsymbol{\hat{S}}$ are the singular values.
The optimal rank-$r$ approximation to $\boldsymbol{A}$, in a least squares sense, is given by the rank-$r$ SVD truncation $\boldsymbol{\tilde{A}}$, [49]:
$\tilde{\boldsymbol{A}}=\tilde{\boldsymbol{U}}\tilde{\boldsymbol{S}}\tilde{\boldsymbol{V}}^{*}=\inf_{\hat{\boldsymbol{A}}:rank(\tilde{\boldsymbol{A}})=r}\lVert\boldsymbol{A}-\hat{\boldsymbol{A}}\lVert_{F}.$ (22)
In the above equation, $\boldsymbol{\tilde{U}}$ and $\boldsymbol{\tilde{V}}$ denote the first $r$ leading columns of $\boldsymbol{U}$ and $\boldsymbol{V}$, and $\boldsymbol{\tilde{S}}$ contains the leading $r\times r$ sub-block of $\boldsymbol{S}$, $\lVert\bullet\lVert_{F}$ is the Frobenius norm. For a given rank $r$, there is no better approximation for $\boldsymbol{A}$, in the $\ell_{2}$ sense, than the truncated SVD approximation, $\tilde{\boldsymbol{A}}$.
In the POD-TANN approach, the generic field $\boldsymbol{a}$ is specialized to the field of Internal Coordinates, $\boldsymbol{\xi}$, and the snapshot matrix $\boldsymbol{A}$ into $\boldsymbol{\Xi}$. Since the POD modes are time-invariant, the time-varying POD coefficients of the low-rank approximation of $\boldsymbol{\xi}$ may be used as a surrogate variable for tracking the irreversibility in the system. Thus, the low-rank SVD decomposition of $\boldsymbol{\Xi}$, $\tilde{\boldsymbol{\Xi}}=\tilde{\boldsymbol{U}}\tilde{\boldsymbol{S}}\tilde{\boldsymbol{V}}^{*}$, is exploited to identify the time-varying reduced order state variables of the macro-system, $\boldsymbol{C}\equiv\boldsymbol{Z}$:
$\boldsymbol{\Xi}\approx\tilde{\boldsymbol{\Xi}}=\tilde{\boldsymbol{U}}\tilde{\boldsymbol{S}}\tilde{\boldsymbol{V}}^{*}=\boldsymbol{\Phi}\boldsymbol{Z}\to\begin{cases}\boldsymbol{\Phi}=\tilde{\boldsymbol{U}},\\ \boldsymbol{Z}=\tilde{\boldsymbol{S}}\tilde{\boldsymbol{V}}^{*}=\tilde{\boldsymbol{U}}^{*}\tilde{\boldsymbol{\Xi}}.\end{cases}$ (23)
The matrix $\boldsymbol{Z}$ collects the time-varying POD coefficients describing the influence of each POD mode at the considered instant of time. The columns of $\boldsymbol{Z}$, $\boldsymbol{\mathcal{Z}}$ can thus be interpreted as the ISV vectors at the subsequent time instants.